# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ E.1 Nonnegative Linear Inverse Problems
E.2 $\ell_{p}$ Loss Problem
Figure 6: The $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$. The results are reported as mean $\pm$ standard deviation.
@@ E.2 $\ell_{p}$ Loss Problem
ShortFW 2.4697e+01 $\pm$ 1.2e+00 4.7621e+01 $\pm$ 8.6e-01 9.2352e-02 $\pm$ 3.7e-03
OpenFW 1.6961e-02 $\pm$ 9.9e-04 4.$0$$\pm$ 1.0e-01 8.7292e-02 $\pm$ 3.5e-03
We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. We generated $A$ from an i.i.d. normal distribution and normalized it so that $\|a_{i}\|=1$. We also generated $\tilde{x}$ from an i.i.d. normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$ (random seed 1234). The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(n,m)=(100,100)$ and $p=1.1$, Figure 6 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations. Table 3 also shows the average performance over 20 different instances. BregFW outperformed the other algorithms. Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
E.3 Phase Retrieval
@@
Figure 7: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via AFW.
Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$. The results are reported as mean $\pm$ standard deviation.
@@
OpenFW 5.2496e-10 $\pm$ 3.9e-11 4.3802e+00 $\pm$ 5.5e-02 6.0353e+01 $\pm$ 1.6e+00
Table 5: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$. The results are reported as mean $\pm$ standard deviation.
@@
OpenFW 9.3513e-11 $\pm$ 4.9e-12 2.6628e+00 $\pm$ 1.7e-02 7.1916e+01 $\pm$ 1.3e+00
Table 6: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms. The results are reported as mean $\pm$ standard deviation.
@@
BregAFW 4.3430e+00 $\pm$ 1.7e+00 9.7578e-01 $\pm$ 9.1e-01 2.3336e-01 $\pm$ 6.4e-02
EucAFW 4.3968e+00 $\pm$ 1.7e+00 1.2212e+00 $\pm$ 2.3e+00 3.9421e-01 $\pm$ 1.6e-01
ShortAFW 4.4092e+00 $\pm$ 1.7e+00 2.5978e-02 $\pm$ 7.3e-03 7.3540e-02 $\pm$ 3.5e-02
OpenAFW 4.4094e+00 $\pm$ 1.7e+00 7.1992e-02 $\pm$ 4.9e-02 5.8998e-02 $\pm$ 2.4e-03
E.4 Low-Rank Minimization
@@
where $P\subset\mathbb{R}^{n\times r}$. We assume that $r\leq n$. This problem is known as low-rank minimization [Dragomir2021-rv]. In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $r\in\mathbb{R}_{+}$ for the low-rank assumption.
@@
There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Dragomir2021-rv]. Additionally, $f(X)$ is weakly convex on any compact set due to Proposition B.1, which follows from the twice continuous differentiability of $f$.
We also compared BregFW with EucFW, ShortFW, and OpenFW. The parameter settings are the same as those in the previous subsection. We generated $X^{*}$ from an i.i.d. uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$ (random seed 42). The initial point $X_{0}$ was generated from an i.i.d. uniform distribution in $[0,1]$. We set $\xi=10\lambda_{\max}(M)$ for $P$. Figure 8 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration. Table 7 presents the average performance over 20 different instances for $(n,r)=(1000,20)$. BregFW performs slightly better than EucFW. OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable. ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
Figure 8: Low-rank minimization for $(n,r)=(1000,20)$.
Table 7: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW 5.6519e-01 $\pm$ 4.2e-02 1.4727e+01 $\pm$ 1.0e+00 1.2602e+01 $\pm$ 4.5e-01
EucFW 5.7765e-01 $\pm$ 4.6e-02 1.7237e+01 $\pm$ 2.0e+00 1.2280e+01 $\pm$ 3.8e-01
ShortFW 2.7101e+08 $\pm$ 1.9e+06 2.1694e+09 $\pm$ 1.5e+07 4.1379e+00 $\pm$ 7.9e-02
OpenFW 1.9782e+00 $\pm$ 3.6e-01 2.1212e+03 $\pm$ 3.1e+02 4.1770e+00 $\pm$ 1.7e-01
E.5 Nonnegative Matrix Factorization
Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$. NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e., $\min_{(W,H)\in P}f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2}$, where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$. The objective function $f$ is weakly convex over $P$ due to Proposition B.1. The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable [Mukkamala2019-mk] with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
[…]
@@ E.5 Nonnegative Matrix Factorization
Table 8: average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$. The results are reported as mean $\pm$ standard deviation.

### v4

@@ F.1 Nonnegative Linear Inverse Problems
F.2 $\ell_{p}$ Loss Problem
Figure 6: Log-log plot of primal and FW gaps on the $\ell_{p}$ loss problem for $(m,n)=(1000,100)$.
Table 3: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$. The results are reported as mean $\pm$ standard deviation.
@@ F.2 $\ell_{p}$ Loss Problem
ShortFW 2.4697e+01 $\pm$ 1.2e+$0$ 4.7621e+01 $\pm$ 8.6e-01 9.2352e-02 $\pm$ 3.7e-03
OpenFW 1.6961e-02 $\pm$ 9.9e-04 4.0077e+$0$ $\pm$ 1.0e-01 8.7292e-02 $\pm$ 3.5e-03
We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. We generated $A$ from an i.i.d. normal distribution and normalized it so that $\|a_{i}\|=1$. We also generated $\tilde{x}$ from an i.i.d. normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$ (random seed 1234). The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(n,m)=(100,100)$ and $p=1.1$, Figure 6 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations. Table 3 also shows the average performance over 20 different instances. BregFW outperformed the other algorithms. Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gaps.
F.3 Phase Retrieval
@@
Figure 7: Log-log plot of primal and FW gaps on phase retrieval for $(m,n)=(200,200)$ and $K=110$ via AFW.
Table 4: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$. The results are reported as mean $\pm$ standard deviation.
@@
OpenFW 5.2496e-10 $\pm$ 3.9e-11 4.3802e+$0$ $\pm$ 5.5e-02 6.0353e+01 $\pm$ 1.6e+$0$
Table 5: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$. The results are reported as mean $\pm$ standard deviation.
@@
OpenFW 9.3513e-11 $\pm$ 4.9e-12 2.6628e+$0$ $\pm$ 1.7e-02 7.1916e+01 $\pm$ 1.3e+$0$
Table 6: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms. The results are reported as mean $\pm$ standard deviation.
@@
BregAFW 4.3430e+$0$ $\pm$ 1.7e+$0$ 9.7578e-01 $\pm$ 9.1e-01 2.3336e-01 $\pm$ 6.4e-02
EucAFW 4.3968e+$0$ $\pm$ 1.7e+$0$ 1.2212e+$0$ $\pm$ 2.3e+$0$ 3.9421e-01 $\pm$ 1.6e-01
ShortAFW 4.4092e+$0$ $\pm$ 1.7e+$0$ 2.5978e-02 $\pm$ 7.3e-03 7.3540e-02 $\pm$ 3.5e-02
OpenAFW 4.4094e+$0$ $\pm$ 1.7e+$0$ 7.1992e-02 $\pm$ 4.9e-02 5.8998e-02 $\pm$ 2.4e-03
F.4 Low-Rank Minimization
@@
where $P\subset\mathbb{R}^{n\times r}$. We assume that $r\leq n$. This problem is known as low-rank minimization (Dragomir et al., 2021). In this paper, we define $P=\{X\in\mathbb{R}^{n\times r}\,|\,\|X\|_{*}\leq\xi\}$, where $\|\cdot\|_{*}$ denotes the nuclear norm and $\xi\in\mathbb{R}_{+}$ for the low-rank assumption.
@@
There exists a constant $L$ such that the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ (Dragomir et al., 2021). Additionally, $f(X)$ is weakly convex on any compact set due to Proposition C.1, which follows from the twice continuous differentiability of $f$.
We also compared BregFW with EucFW, ShortFW, and OpenFW. The parameter settings are the same as those in the previous subsection. We generated $X^{*}$ from an i.i.d. uniform distribution in $[0,1]$, normalized each column of $X^{*}$, and set $M=X^{*}(X^{*})^{\mathsf{T}}$ (random seed 42). The initial point $X_{0}$ was generated from an i.i.d. uniform distribution in $[0,1]$. We set $\xi=10\lambda_{\max}(M)$ for $P$. Figure 8 shows the primal and FW gaps per iteration and gaps per second for $(n,r)=(1000,20)$ up to the 1000th iteration. Table 7 presents the average performance over 20 different instances for $(n,r)=(1000,20)$. BregFW performed slightly better than EucFW. OpenFW also performed as fast as BregFW and EucFW, but its performance was unstable. ShortFW did not converge due to the lack of Lipschitz continuity of $\nabla f$.
Figure 8: Log-log plot of primal and FW gaps on low-rank minimization for $(n,r)=(1000,20)$.
Table 7: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of low-rank minimization for $(n,r)=(1000,20)$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW 5.6519e-01 $\pm$ 4.2e-02 1.4727e+01 $\pm$ 1.0e+$0$ 1.2602e+01 $\pm$ 4.5e-01
EucFW 5.7765e-01 $\pm$ 4.6e-02 1.7237e+01 $\pm$ 2.0e+$0$ 1.2280e+01 $\pm$ 3.8e-01
ShortFW 2.7101e+08 $\pm$ 1.9e+06 2.1694e+09 $\pm$ 1.5e+07 4.1379e+$0$ $\pm$ 7.9e-02
OpenFW 1.9782e+$0$ $\pm$ 3.6e-01 2.1212e+03 $\pm$ 3.1e+02 4.1770e+$0$ $\pm$ 1.7e-01
F.5 Nonnegative Matrix Factorization
Given a nonnegative matrix $V\in\mathbb{R}_{+}^{m\times n}$, nonnegative matrix factorization (NMF) aims to find nonnegative matrices $W\in\mathbb{R}_{+}^{m\times r}$ and $H\in\mathbb{R}_{+}^{r\times n}$ such that $V\simeq WH$. NMF can be formulated as a minimization problem of the loss function that measures the difference between $V$ and $WH$, i.e., $\min_{(W,H)\in P}f(W,H):=\frac{1}{2}\|WH-V\|_{F}^{2}$, where $P$ is a compact convex subset of $\mathbb{R}_{+}^{m\times r}\times\mathbb{R}_{+}^{r\times n}$. The objective function $f$ is weakly convex over $P$ due to Proposition C.1. The gradient $\nabla f$ is not Lipschitz continuous, while $(f,\phi)$ is smooth adaptable (Mukkamala and Ochs, 2019) with $\phi(W,H)=\frac{1}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{1}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$.
We used a box constraint $P=\{(W,H)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\,|\,0\leq W_{lj}\leq 3,0\leq H_{lj}\leq 1\}$. We compared BregFW with EucFW because ShortFW and OpenFW stopped at the 2nd iteration. We generated $W^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized each column of $W^{*}$ (random seed 42). We also generated $H^{*}$ from an i.i.d. Dirichlet distribution. The initial point $(W_{0},H_{0})$ was generated from an i.i.d. uniform distribution in $[0,1]$. Figure 10 shows the primal and FW gaps for $(m,n,r)=(100,5000,20)$ up to the 5000th iteration. Table 8 shows the average performance over 20 different instances for $(m,n,r)=(100,5000,20)$ up to the 5000th iteration. BregFW is slightly better than EucFW in terms of the primal gap, while the FW gap for EucFW is smaller than that for BregFW.
Furthermore, we consider real-world data using the MovieLens 100K Dataset. We set $P=\{(W,H)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\,|\,0\leq W_{lj}\leq 5.0,\|H\|_{*}\leq\xi\}$. For $\phi(W,H)=\frac{3}{4}(\|W\|_{F}^{2}+\|H\|_{F}^{2})^{2}+\frac{\|V\|_{F}}{2}(\|W\|_{F}^{2}+\|H\|_{F}^{2})$ by (Mukkamala and Ochs, 2019, Proposition 2.1), $(f,\phi)$ is also $L$-smooth adaptable. Figure 10 shows the primal and FW gaps up to the 1000th iteration with $\xi=10\sqrt{\lambda_{\max}(VV^{\mathsf{T}})}$, where $\lambda_{\max}(\cdot)$ is the largest eigenvalue. The primal gap of BregFW is 6.041269e+05, and that of EucFW is 6.041272e+05. In this setting, the primal and FW gaps of BregFW are slightly better than those of EucFW.
@@ F.5 Nonnegative Matrix Factorization
Table 8: Average values of the primal gap, FW gap, and computational time (s) over 20 different instances of NMF for $(m,n,r)=(100,5000,20)$. The results are reported as mean $\pm$ standard deviation.